OpenStudy (anonymous):

check my answer police can estimate the speed of a vehicle before the brakes are applied using the formula 0.75d=s^2/30.25, where s is the speed in miles per hour and d is the length of the vehicle's skid marks. what was the approximate speed of a vehicle that left skid marks measuring 140 feet? A. about 10 miles per hour <-- my answer B. about 75 miles per hour C. about 15 miles per hour D. about 56 miles per hour

OpenStudy (anonymous):

I got D. .75(140)=s^2/30.25 Distribute: 105=s^2/30.25 Multiply by 30.25: 3176.25=s^2 Find the square root of 3176.25